Calculate pI of the following molecules: 6 pts Alanine b. Arginine   c. Ile-Gln

  1. Calculate pI of the following molecules: 6 pts

  1. Alanine b. Arginine   c. Ile-Gln
Answer #1

pI is that pH at which molecule exist in neutral state.

a) For neutral Amino acid, alanine :

pI is the average of pKa of both COOH and NH3+ groups.

pI = (9.69+2.34)/2 = 6.01

b) For basic Amino acid, it is average of both basic groups.

pI = (9.04+12.48)/2 = 10.76

c) For peptide it is the average of pKa of groups before and after neutral form.

pI = (2.36+9.13)/2 = 5.74